Wendy's wonders how to price its new chili salad. It test-markets five different prices in the cities listed below, with the listed resultant sales (in thousands of dollars). Manually calculate correlation, and the regression model. Show your work.

City Sales($000s) Price($) Price^2 Sales^2 Price*Sales
Rochester 75 0.99 0.98 5625.00 74.3
Ottumwa 70 1.29 1.66 4900.00 90.3
Seattle 45 1.49 2.22 2025.00 67.1
Raleigh 33 1.89 3.57 1089.00 62.4
Denair 42 2.19 4.80 1764.00 92.0
sum 265 7.85 13.23 15403.00 386.0
mean 53 1.57 2.65 3080.60 77.2
st.dev. 18.4 0.5 1.53 2037.06 13.4
Answer #1

X Values
∑ = 265
Mean = 53
∑(X - Mx)2 = SSx = 1358

Y Values
∑ = 7.85
Mean = 1.57
∑(Y - My)2 = SSy = 0.908

X and Y Combined
N = 5
∑(X - Mx)(Y - My) = -30.1

R Calculation
r = ∑((X - My)(Y - Mx)) / √((SSx)(SSy))

r = -30.1 / √((1358)(0.908)) = -0.8572

Sum of X = 265
Sum of Y = 7.85
Mean X = 53
Mean Y = 1.57
Sum of squares (SSX) = 1358
Sum of products (SP) = -30.1

Regression Equation = ŷ = bX + a

b = SP/SSX = -30.1/1358 = -0.0222

a = MY - bMX = 1.57 - (-0.02*53) = 2.7447

ŷ = -0.0222X + 2.7447